OpenStudy (anonymous):

Find a third–degree polynomial function such that f(0) = –12 and whose zeros are 1, 2, and 3. Using complete sentences, explain how you found it.

OpenStudy (anonymous):

since the zeros are given, then you know what the polynomial is in factored form... can you give the factors?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a zeros is where the function crosses the x-axis... and since the zeros of the polynomial is given as x=1, 2, 3, then your polynomial will hold true for 0 = a(x-1)(x-2)(x-3)

OpenStudy (anonymous):

can you multiply out (x-1)(x-2)(x-3) for me?

OpenStudy (anonymous):

x^3 - 6x^2 + 11x -6

OpenStudy (anonymous):

ok.. so the equation will look like this: \(\large y=a(x^3 - 6x^2 + 11x - 6) \) now you have to meet the condition f(0) = -12

OpenStudy (anonymous):

distribute the "a" so you'll have \(\large y=ax^3-6ax^2+11ax-6a \) when x=0, y=-12. you must figure out what "a" is for this to be true.